OpenStudy (anonymous):

if i get a denominator with one of its factor (x^3 - 2) It's numerator in partial fraction decomposition should be Ax^2 + Bx + C am I right? if not, what's right?

OpenStudy (anonymous):

hmm can you explain your question a little more

OpenStudy (anonymous):

I am working with a technique of integration called partial fraction decomposition. One of the factors of the denominator is (x^3 - 2). I'm wondering if It's Numerator should be Ax^2 + Bx + C or not. Help me please! thanks :D

OpenStudy (anonymous):

\[(x^3 - 2) = (x - 2^{\frac{1}{3}})(x^2 + 4^{\frac{1}{3}} + x*2^{\frac{1}{3}})\]